TPC weekly

2021-04-21 15:00
America/New York
Wednesday, 21 April 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC meeting April21 minutes

 

Present:

Alexei,Babu, Flemming,Gene,Hongwei,Yuri

 

TPC running well. The pizo driver for the laser failed , no power,was adjusted during quick ccess, and then rebooted today during access. Lost about one laser run during. Will investigate how remote reboot can be done. Such failure has happened 1-2 per run.

 

Flemming present a few more observation on the single bad pad hot due to values this seem to have pedestal subtracted incorrectly. This is intermittent and usually, but not always clears on next run. Also it only 5-10% of events with this failure. See  https://drupal.star.bnl.gov/STAR/blog/videbaks/tpc-qa-2021
for more details

 

The 19.6 GeV calibration sample is complete. Includes dynamic space charge, and event by event t0 correction. EPD had time offsets calibrated/changed at some point in time during the 19.6 GeV data sets. This change is now in DB.  The 14.5 GeV  should be done by the weekend. Babu will start looking at the run 20 soon.

 

The sample(s) can now be checked for BTOF calibrations, and dE/dx.

If the mass square difference seen in east persists with the DbTpcMaker that was updated, there will be additional work to identify this issue. The TOF group will not sign off with a mass split of approx. 1%.